OpenStudy (anonymous):

The population of a local species of dragonfly can be found using an infinite geometric series where a1 = 36 and the common ratio is one half. Write the sum in sigma notation and calculate the sum that will be the upper limit of this population

OpenStudy (anonymous):

Here are the answer choices if you need them to look at

hartnn (hartnn):

The general formula for n'th term of Geometric sequence is \(\Large a_i = a_1 r^{i-1}\) here, a1 = 36 and r = 1/2 just plug in!

hartnn (hartnn):

and for the 2nd part, use this sum formula \(\Large S_\infty = \dfrac{a_1}{r-1}\)

hartnn (hartnn):

sorry, that should be \(\Large S_\infty = \dfrac{a_1}{1-r}\)

OpenStudy (anonymous):

\[a _{i}= 36 ^{1/2-1}\] is what you meant?

hartnn (hartnn):